# latex‘h’ float specifier changed to ‘ht’

### v1

@@ 5.1. Multi-Player Game among Builders $\left(M\geq 3\right)$
    Each player’s ability to extract MEV without the order being auctioned in the OFA is $k_{1,i}\coloneqq\frac{\bar{v}_{i}}{\bar{f}_{i}}$imes the MEV obtained from the auctioned order, and this ratio $k_{1,i}$emains constant across all three players, i.e., $k_{1,1}=k_{1,2}=k_{1,3}$
@@
    The ratio $k_{1,i}$aries among the three players, such that $k_{1,1}<k_{1,2}<k_{1,3}$
The results for three players are presented in Table 2 and Table 3.
@@
$\mathbf{E}\left(\pi_{3}\right)$ 13.00 20.34 35.14 57.40 72.25
@@
$\mathbf{E}\left(\pi_{3}\right)$ 13.74 27.48 41.22 68.71 109.93
@@ 6. Conclusions
In this paper, we examine the interaction between the order flow auction and the block-building auction, formulating the problem within a general multiplayer framework. For the case of two players, we establish the existence and uniqueness of the Nash equilibrium, while for more than two players, we conduct simulations. Both analyses suggest a tendency toward centralization in the builder space. In contrast, we find that validators’ stake shares follow a martingale process, and both theoretical and numerical results indicate that centralization is unlikely to emerge in the validator space.
@@
Appendix A Code for the Problem 3.8
@@
$\text{Var}\left(\omega_{j,t+1}\right)=a_{t}\omega_{j,0}\left(1-\omega_{j,0}\right),$ (B.1)
@@
We prove by induction, following the approach of Rosu and Saleh [18]. Since $\mathbf{E}_{t}\left(\omega_{j,t+1}\right)=\omega_{j,t}$, the following equation holds for all $t\geq 0$:
@@
where the last inequality follows from the fact that $S_{i}=S_{0}+\sum_{j=1}^{i}R_{i}-i\alpha$. Since the function $x\to\left(\frac{R_{\max}}{S_{0}+x\left(R_{\min}-\alpha\right)}\right)^{2}$ is decreasing on $\mathbb{R}_{+}$ due to $R_{\max}>0,S_{0}>0$, and $R_{\min}-\alpha>0$ (by Assumption 2.10), using the sum-integral trick as in  Rosu and Saleh [18, Lemma A.4], we obtain:
[…]

### v2

@@ 5.1. Multi-Player Game among Builders $\left(M\geq 3\right)$
    Each player’s ability to extract MEV without the order being auctioned in the OFA is $k_{1,i}\coloneqq\frac{\bar{v}_{i}}{\bar{f}_{i}}$ times the MEV obtained from the auctioned order, and this ratio $k_{1,i}$ remains constant across all three players, i.e., $k_{1,1}=k_{1,2}=k_{1,3}$.
@@
    The ratio $k_{1,i}$ varies among the three players, such that $k_{1,1}<k_{1,2}<k_{1,3}$.
The results for four players are presented in Table 4 and Table 5 in Appendix C.
@@
$\mathbf{E}\left(\pi_{3}\right)$ 4.37 6.94 12.13 19.95 25.18
@@
$\mathbf{E}\left(\pi_{3}\right)$ 0.35 0.71 1.06 1.77 2.83
@@ 6. Conclusions
In this study, we examine the interactions between the order flow auction and the block-building auction, formulating the problem as a multiplayer game. We establish the existence of a Nash equilibrium in the general setting and prove its uniqueness in the two-player case, along with deriving the closed-form solution. Furthermore, we show that validators’ stake shares follow a martingale process. We also conduct numerical simulations to complement our theoretical analysis. Both theoretical and numerical results suggest a tendency toward centralization in the builder space, which, in contrast, is highly unlikely in the validator space.
@@
Appendix A Code for Problem 3.8
@@
$\text{Var}\left(\omega_{j,t}\right)=a_{t}\omega_{j,0}\left(1-\omega_{j,0}\right),$ (B.1)
@@
We prove by induction, following the approach of rosu2021market. Since $\mathbf{E}_{t}\left(\omega_{j,t+1}\right)=\omega_{j,t}$, the following equation holds for all $t\geq 0$:
@@
where the last inequality follows from the fact that $S_{i}=S_{0}+\sum_{j=1}^{i}R_{i}-i\alpha$. Since the function $x\to\left(\frac{R_{\max}}{S_{0}+x\left(R_{\min}-\alpha\right)}\right)^{2}$ is decreasing on $\mathbb{R}_{+}$ due to $R_{\max}>0,S_{0}>0$, and $R_{\min}-\alpha>0$ (by Assumption 2.8), using the sum-integral trick as in  rosu2021market, we obtain:
$a_{t}\leq\int_{0}^{\infty}\left(\frac{R_{\max}}{S_{0}+x\left(R_{\min}-\alpha\right)}\right)^{2}dx=\frac{R_{\max}^{2}}{S_{0}\left(R_{\min}-\alpha\right)}.$ (B.6)
@@ Appendix C Additional Results for Numerical Experiments
$\bar{f}_{i}/\bar{v}_{i}$ 2 3 5 8 10
$h_{1}$ 154.60 209.39 319.68 485.63 596.37
$\mathbf{E}\left(\pi_{1}\right)$ 94.89 133.59 211.51 328.74 406.96
$h_{2}$ 96.45 134.13 209.45 322.38 397.67
$\mathbf{E}\left(\pi_{2}\right)$ 30.28 45.02 74.48 118.65 148.09
$h_{3}$ 63.84 90.77 144.32 224.47 277.87
$\mathbf{E}\left(\pi_{3}\right)$ 12.40 19.03 32.30 52.21 65.49
$h_{4}$ 31.16 45.46 73.91 116.47 144.82
$\mathbf{E}\left(\pi_{4}\right)$ 2.84 4.49 7.82 12.83 16.18
Table 4. This table presents the equilibrium outcomes for four players under the assumption that each player $i\in\left[4\right]$ has a common ratio $\bar{f}_{i}/\bar{v}_{i}$, where the ratio is set to $2,3,5,8$, or $10$. The players’ $\bar{v}_{i}$ follow the ratio $\bar{v}_{1}:\bar{v}_{2}:\bar{v}_{3}:\bar{v}_{4}=5:3:2:1$, with $\bar{v}_{4}=30$.
$\bar{v}_{4}$ 10 20 30 50 80
$h_{1}$ 13099.66 26199.33 39298.99 65498.31 104797.30
$\mathbf{E}\left(\pi_{1}\right)$ 23844.40 47688.79 71533.19 119221.98 190755.16
$h_{2}$ 6164.17 12328.35 18492.52 30820.87 49313.39
$\mathbf{E}\left(\pi_{2}\right)$ 2687.11 5374.23 8061.34 13435.56 21496.90
$h_{3}$ 976.29 1952.57 2928.86 4881.44 7810.30
$\mathbf{E}\left(\pi_{3}\right)$ 49.30 98.61 147.91 246.52 394.43
$h_{4}$ 49.96 99.93 149.89 249.81 399.70
$\mathbf{E}\left(\pi_{4}\right)$ 0.12 0.25 0.37 0.62 0.99
Table 5. This table presents the equilibrium outcomes for four players under the assumption that $\frac{\bar{f}_{1}}{\bar{v}_{1}}=1000>\frac{\bar{f}_{2}}{\bar{v}_{2}}=500>\frac{\bar{f}_{3}}{\bar{v}_{3}}=100>\frac{\bar{f}_{4}}{\bar{v}_{4}}=10$. The players’ $\bar{v}_{i}$ follow the ratio $\bar{v}_{1}:\bar{v}_{2}:\bar{v}_{3}:\bar{v}_{4}=5:3:2:1$, where $\bar{v}_{4}$ is set to $10,20,30,50$, or $80$.